An electrical mini speaker for home requires 9 V to function. The voltage of the electrical supply is 240 V. What is the turns ratio of the transformer that must be used?
1
Expert's answer
2020-04-30T10:39:15-0400

From the formula for the transformation coefficient

"k=\frac{U_1}{U_2}=\frac{N_1}{N_2}"

the ratio of the number of turns of the transformer is equal

"\frac{N_1}{N_2}=\frac{U_1}{U_2}=\frac{240}{9}=26.67"